Question 20. Cost of painting 4 walls of a room 90 meters x 40 meters at Rs. 10 per sq.m is Rs.13,000. Find the height of the room

Answer by mananth(16946)

cost of painting = 13000
rate /sq. meter = 10
area = 13000/10= 1300 sq. meters
..
let height of the room be h
Area of painting = (2*h*W)+(2*h*L) ( 2 walls + 2 walls)
1300= 2*h*40+2*h*90
1300= 80h+180h
1300=260 h
h= 1300/260
h=5 meters the height of the room